OpenStudy (anonymous):

sinθ=sin((1)/(2))θcos((1)/(2))θ?

OpenStudy (raden):

can u draw ur equation ? too many brackets there :P

geerky42 (geerky42):

\[\sinθ= \sin\left(\dfrac{(1)}{(2)}\right)\theta \cos\left(\dfrac{(1)}{(2)}\right)\theta\] It makes no sense, so no. It's not true.